A 1-kg object moving at 1 m/s undergoes a one-dimensional elastic collision with a 2-kg object initially at rest. What is the speed of the 1-kg object after the collision?


1
Expert's answer
2021-01-22T09:56:49-0500

"m_1 = 1 \;kg \\

u_1 = 1 \;m/s \\

m_2 = 2 \;kg \\

u_2 = 0 \;m/s"

Final speed of the object (1):

"v_1 = \frac{m_1-m_2}{m_1+m_2} \times u_1 + \frac{2m_2}{m_1+m_2} \times u_2 \\

= \frac{1-2}{3} \times 1 + \frac{2 \times 2}{3} \times 0 \\

= -\frac{1}{3} \\

= -0.33 \;m/s"

Negative sign shows direction.

Answer: 0.33 m/s